if philip's utility function is... Question: If philip's utility function is U=2q1^0.5+q2, what are his demand functions for two goods? if philip's utility function is U=2q1^0.5+q2, what are his demand functions for two goods
1
Expert's answer
2021-08-03T13:27:53-0400

given utility function "U(x,y)=2q_1^{0.5}+q_2"

"MRS=\frac{MU_{q_1}}{MU_{q_2}}"


"=\frac{\frac{\delta U(q_1,q_2)}{\delta q_1}}{\frac{\delta U(q_1,q_2)}{\delta q_2}}"


"=\frac{0.5q_1^{-0.5}}{1}"


"=\frac{1}{2q_1^{0.5}}"

budget line "p_1q_1+p_2q_2=M"

slope of budget line is "\frac{p_1}{p_2}"

The optimal consumption is where MRS equal slope of budget line. it is shown below: "\frac{1}{2q_1^{0.5}}=\frac{p_1}{p_2}"


"\frac{p_2}{2p_1}=q_1^{0.5}"

squaring both sides

"(\frac{p_2}{2p_1})^2=q_1"


substitute this value of q1 into budget line

"p_1q_1+p_2q_2=M\\p_1(\frac{p_2}{2p_2})^2+p_2q_2=M\\\frac{p_2^2}{4p_1}+p_2q_2=M\\P_2Q_2=M-\frac{p_2^2}{4p_1}\\q_2=\frac{M}{p_2}-\frac{p_2}{4p_1}"

Thus , the demand function of q1 "(\frac{p_2}{2p_1})^2" is and q2 is "\frac{M}{p_2}-\frac{p_2}{4p_1}"